2.104 An insurance company offers its policyholders a number of different premium payment options. For a randomly selected policyholder, let X be the number of months between successive payments. The cumulative distribution function of X is 0, if ac 1, 0.4 if 1 z 3, F (r) 0.6 if 3 S z 5, 0.8 if 5 r 7, 1.0 if ac 7 (a) What is the probability mass function of X? (b) Compute P (4 <X S 7).
Answer #1

(a) P(X=1) = 0.4 P(X=3)=0.6-0.4-0.2 P (X = 5) = 0.8-0.6-0.2 P(X=7)-1-0.8 = 0.2 Thus the PMF of X is, P(x 0.4 0.2 0.2 0.2 (b)